# Is Table Retrieval a Solved Problem? Join-Aware Multi-Table Retrieval

### v1

@@ Abstract
Join-Aware Multi-Table Retrieval
@@ 1 Introduction
Figure 1: Previous work make simplifying assumptions for table retrieval. Left block assumes there exists a single table that is sufficient to answer the question. Middle block assumes necessary joins can be recovered from query decomposition. However, in practice, the question may be more complicated and the join plan may not be discerned from the user query itself (right block). Therefore it is important to identify the join relationship while conducting table retrieval.
@@ 3.1 Query-table relevance
A standard way to compute query-table relevance is to compute a similarity score through a bi-encoder model architecture Chen et al. (2020); Huang et al. (2022); Herzig et al. (2021). Such a similarity score can capture coarse-grained relevance. However, real queries can be complicated (e.g., multi-hop questions). The relevant information to answer the query may be distributed across multiple tables, and each table that should be used to answer the query may be only mapped to a part of the query. In our running example, a Card table may only tell us "who own credit cards" but not "who have loans".
@@
3.2 Table Re-ranking
As discussed above, we prefer tables with high table-query relevance and table-table relevance.
For table-query relevance, we consider both coarse-grained and fine-grained relevance introduced in Section 3.1, so that we can find the tables that are most relevant to all the information that is required to answer the query. To ensure the coverage of different information, we further complement the relevance score with a coverage measurement, which will be described in detail in Section 3.2.1.
For table-table relevance, we consider the compatibility among the returned tables, i.e., whether selected tables can join each other. Joinability helps us ensure that information from the selected tables is compatible — the information is about the same group of entities or objects, so we can use all of them to answer the query correctly. Therefore, while selecting tables that can maximize the compatibility scores, we further constrain the selected tables to be connected, which will be discussed in Section 3.2.2.
@@ 3.2.1 Sub-query Coverage
The main purpose of measuring fine-grained table-query relevance is to find different columns (may or may not come from the same table) that can map to different parts of the query, so that all information required to answer the query can be covered by different columns. Although we have $\sum_{q,i,k}r_{qik}d_{qik}$ in the objective function to maximize the overall relevance between different parts of the query and columns, it may be biased to aligning a sub-query (weakly) to many columns to maximize the term rather than aligning a sub-query strongly to one column.
@@ 3.3 Table-Table Relationship Inference
Compatibility between two tables can be represented as the most compatible pair of columns between the two tables. To compute the compatibility between two columns, we consider two aspects: column relevance and the likelihood of satisfying the key foreign-key constraint.
@@
To summarize, compatibility score $\omega_{ij}^{kl}$ comprises two components: (1) column relevance that includes both instance similarity $j$ and schema similarity $e$ (2) likelihood of satisfying key-foreign key constraint, which is the maximum of the uniqueness $u$ of two columns. Specifically, $\omega_{ij}^{kl}=(e(c_{k},c_{l})+j(c_{k},c_{l}))*\max\{u(c_{k}),u(c_{l})\}$.
@@ 4 Evaluation
In this section, we evaluate the baselines and our re-ranking mechanism on datasets whose questions require multiple tables to answer. The goal is to answer the following two research questions: (1) To what extent can our re-ranking mechanism, which jointly considers table-query relevance and table-table relevance, improve the existing table retrieval solutions that consider only standard table-query relevance? (2) To what extent can a better retrieval performance due to our re-ranking mechanism enhance the end-to-end performance of question answering over tables?
@@ 4.1 Experimental Settings
To the best of our knowledge, there are no existing large-scale open-domain question-answering datasets that consider multiple tables [^2] KaggleDBQA Lee et al. (2021) has only 26 queries involving multiple tables, which is not a representative dataset for our task.. Therefore, we use text-to-SQL datasets which naturally consider multiple tables for our evaluation, but under an open-domain question-answering setup. In this case, it is required to first retrieve tables from a table corpus, and then answer the query by generating an SQL expression. Specifically, we run evaluation on two popular text-to-SQL datasets: Spider Yu et al. (2018b) and Bird Li et al. (2023a). In these two datasets, tables are organized by topic, and each topic has its own database (around 5.4 tables per database) and queries that can be answered. To make these datasets suitable for our setting, we aggregate tables from all databases to construct a centralized table corpus for each dataset. The original datasets contain key foreign-key constraints. However, as described in Section 1, such constraints are not always specified and may need to be inferred during test time. Therefore, we perform evaluations on two setups: with and without the gold key foreign-key constraints provided by the original datasets (if gold constraints are provided, we set the compatibility score of the corresponding pairs of columns to 1). Moreover, since we focus on queries that require multiple tables to answer, we omit those that only use one table to construct the gold SQL expression. We also remove queries with incorrect answers through human inspection. We provide examples to explain the removal process in Appendix C. The above procedure leads to 443 queries and 81 tables for Spider, and 1095 queries and 77 tables for Bird.
@@
In the following sections, we discuss the results for table retrieval in Section 4.2 and the end-to-end question answering task in Section 4.3. We also include an ablation study and a discussion of the results when gold key foreign-key constraints are available in Section 4.4.
@@ 5 Conclusion
Table retrieval is critical for open-domain question answering, fact-checking, and, more generally, retrieval-augmented generation that relies on information from tables. However, existing work mostly focuses on single table retrieval without considering the real-life scenarios where tables are typically normalized, and therefore, multiple tables need to be joined to produce sufficient information. In the case they do consider multi-table retrieval, they only consider query-table relevance. In this paper, we proposed a re-ranking method based on mixed-integer programming to select the best set of tables as a result of joint decisions between both relevance and compatibility. Our experiments demonstrated that our approach outperforms the baseline approaches in standard retrieval performances as well as end-to-end evaluation on the question answering downstream task.
@@ 6 Limitations
The main focus of this paper is to encourage the community to consider multiple tables when conducting table retrieval for downstream tasks, such as open-domain question answering. The proposed MIP-based reranking mechanism in this paper takes a first step to consider join relationship over tables during the inference time. However, in a real-life scenario, join is not the only relationship existing among tables. Another common relation is union, i.e., answers to user queries can be distributed into multiple tables horizontally, and only through a union operation over those tables can the model return a complete answer. Detecting unionable tables during retrieval differs from detecting joinable tables, which we did not explore in this paper. Furthermore, how to combine joinable tables and unionable tables to return complete answers is also an interesting future work.

### v2

# Is Table Retrieval a Solved Problem? Exploring Join-Aware Multi-Table Retrieval

@@ Abstract
Exploring Join-Aware Multi-Table Retrieval
@@ 1 Introduction
Figure 1: Previous work makes simplifying assumptions for table retrieval. Left block assumes there exists a single table that is sufficient to answer the question. Middle block assumes necessary joins can be recovered from query decomposition. However, in practice, the question may be more complicated and the join plan may not be discerned from the user query itself (right block). Therefore it is important to identify the join relationship while conducting table retrieval.
@@ 3.1 Query-Table Relevance
A standard way to compute query-table relevance is to compute a similarity score through a bi-encoder model architecture Chen et al. (2020b); Huang et al. (2022); Herzig et al. (2021). Such a similarity score can capture coarse-grained relevance. However, real queries can be complicated (e.g., multi-hop questions). The relevant information to answer the query may be distributed across multiple tables, and each table that should be used to answer the query may be only mapped to a part of the query. In our running example, a Card table may only tell us "who own credit cards" but not "who have loans".
@@
3.2 Table Re-ranking with Table-Table Relevance
As discussed above, we prefer tables with high query-table relevance and table-table relevance.
For query-table relevance, we consider both coarse-grained and fine-grained relevance introduced in Section 3.1, so that we can find the tables that are most relevant to all the information that is required to answer the query. To ensure the coverage of different information, we further complement the relevance score with a coverage measurement, which will be described in detail in Section 3.2.1.
For table-table relevance, we consider the compatibility among the returned tables, i.e., whether selected tables can join with each other. The compatibility (joinability) helps us ensure that the information from the selected tables is compatible — the information is about the same group of entities or objects, so that we can use all of them to answer the query correctly. Therefore, during the process of selecting tables that can maximize the compatibility scores, we further encourage the selected tables to be chained through join, which will be discussed in Section 3.2.2.
@@ 3.2.1 Sub-query Coverage
The main purpose of measuring fine-grained table-query relevance is to find different columns (may or may not come from the same table) that can map to different parts of the query, so that all information required to answer the query can be covered. Although we have $\sum_{q,i,k}r_{qik}d_{qik}$ in the objective function to maximize the overall relevance between different parts of the query and columns, it may be biased to aligning a sub-query (weakly) to many columns to maximize the term rather than aligning a sub-query strongly to one column.
@@ 3.3 Table-Table Relationship Inference
The compatibility between two tables can be represented as the most joinable pair of columns between the two tables. To compute the joinablity between two columns, we consider two aspects: column relevance and the likelihood of satisfying the key foreign-key constraint.
@@
To summarize, joinablity score $\omega_{ij}^{kl}$ comprises two components: (1) column relevance that includes both instance similarity $j$ and schema similarity $e$ (2) likelihood of satisfying key-foreign key constraint, which is the maximum of the uniqueness $u$ of two columns. Specifically, $\omega_{ij}^{kl}=(e(c_{k},c_{l})+j(c_{k},c_{l}))*\max\{u(c_{k}),u(c_{l})\}$.
@@ 4 Evaluation
In this section, we evaluate the baselines and our re-ranking method on datasets whose questions require multiple tables to answer. The goal is to answer the following two research questions: (1) To what extent can our re-ranking method, which jointly considers query-table relevance and table-table relevance, improve the existing table retrieval solutions that consider only standard query-table relevance? (2) To what extent can a better retrieval performance due to our re-ranking method enhance the end-to-end performance of question answering over tables?
@@ 4.1 Experimental Settings
To the best of our knowledge, there are no existing large-scale open-domain question-answering datasets that consider multiple tables [^2] KaggleDBQA Lee et al. (2021) has only 26 queries involving multiple tables, which is not a representative dataset for our task.. Therefore, we use text-to-SQL datasets which naturally consider multiple tables for our evaluation, but under an open-domain question-answering setup. In this case, it is required to first retrieve tables from a table corpus, then answer the query by generating an SQL expression. Specifically, we run evaluation on two popular text-to-SQL datasets: Spider Yu et al. (2018b) and Bird Li et al. (2023a). In these two datasets, tables are organized by topic, and each topic has its own database (around 5.4 tables per database) and queries that can be answered. To make these datasets suitable for our setting, we aggregate tables from all databases to construct a centralized table corpus for each dataset. The original datasets contain key foreign-key constraints. However, as described in Section 1, such constraints are not always specified and may need to be inferred during test time. Therefore, we perform evaluations on two setups: with and without the gold key foreign-key constraints provided by the original datasets (if gold constraints are provided, we set the compatibility score of the corresponding pairs of columns to 1). Moreover, since we focus on queries that require multiple tables to answer, we omit those that only use one table to construct the gold SQL expression. We also remove queries with incorrect answers through human inspection. We provide examples to explain the removal process in Appendix C. The above procedure leads to 443 queries and 81 tables for Spider, and 1095 queries and 77 tables for Bird.
@@
In the following sections, we discuss the results for table retrieval in Section 4.2 and the end-to-end question answering task in Section 4.3. We also include an ablation study as well as a discussion of the results when gold key foreign-key constraints are available and results under different datasets and numbers of tables in Section 4.4.
@@ 4.4 Discussion
Queries with 2 tables Queries with $\geq 3$ tables Spider Bird Spider Bird P R F1 P R F1 P R F1 P R F1 Table retrieval performances (Top-5) DTR JAR-D 38.8 97 55.4 36.1 90.3 51.6 58.9 94.7 72.5 50.8 82.7 62.8 JAR-F 39.2 98 56 36.5 91.3 52.2 59.6 95.9 73.4 52 84.6 64.3 Diff. 0.4 1.0 0.6 0.4 1.0 0.6 0.7 1.2 0.9 1.2 1.9 1.5 Contriever JAR-D 37.9 94.8 54.2 35.9 89.7 51.2 52.3 83.6 64.2 49.8 80.7 61.5 JAR-F 38 94.9 54.2 36.6 91.6 52.3 53 85.1 65.2 52 84.5 64.3 Diff. 0.1 0.1 0.0. 0.7 1.9 1.1 0.7 1.5 1.0 2.2 3.8 2.8
Table 4: Our approach is the most effective on more challenging and realistic queries involving more tables.
Retrieval performances under different datasets and numbers of tables.
Table 4 shows that table-table relevance contributes (as measured by the difference between the retrieval performances of JAR-F and JAR-D) more to retrieval performances for (1) queries involving more tables and (2) Bird compared to Spider. Since our focus is to improve table-related downstream tasks involving multiple tables, the fact that table-table relevance provides the most benefits with a higher number of tables should be expected. Moreover, we find that queries in the Bird dataset are more challenging and realistic because they involve more bridging tables (e.g., the Disp table as mentioned in Figure 1) that are less directly relevant to the user question. They also contain more confusing tables (tables with similar schema, such as Client (client_id) and Customers (CustomerID)). These features resemble what exists in real life and table-table relevance can address these more challenging issues, demonstrating its effectiveness.
5 Related Work
One line of related work is text-to-SQL Zhong et al. (2017); Yu et al. (2018b). Thanks to recent advances in LLM, SQL queries (including tables) corresponding to a natural language question or query can be generated through prompting engineering Nan et al. (2023); Liu et al. (2023) or fine-tuning Gao et al. (2024); Wang et al. (2024) with LLMs. Although, in this paper, we evaluate our method on the aggregated version of text-to-SQL datasets, we target a general multi-table retrieval problem that requires the consideration of multiple (joinable) tables. In other words, we do not limit our work to specific downstream tasks (e.g., text-to-SQL for a given database).
To the best of our knowledge, we are the first work to propose the problem of join-aware multi-table retrieval for complex questions, and our paper encourages considering table relationships for retrieval during inference time. One related direction from the literature is fusion retrieval Chen et al. (2020a). The basic idea of fusion retrieval is based on an “early fusion” strategy to group relevant heterogeneous data (including text and tables) before retrieval. The early fusion process aims to fuse a table segment and relevant passages into a group. It is implemented by linking entities mentioned in a table segment to the appropriate passages, similar to traditional document expansion based on individual entities. Although this method can also capture the relationship between table and text, it cannot capture key foreign-key relationships among multiple tables. Therefore, when answering aggregation questions, false-positive tables are very likely to be introduced due to the expansion through individual entities. Furthermore, tables do not have to be connected by columns of entities, and the foreign key column can be just as simple as an id column. Last but not least, real questions may require more than 2 tables to answer, and therefore bridging tables may be needed to connect tables. Sometimes, the bridging tables can also cross multiple hops. In this case, tables introduced by an "early fusion" strategy will grow exponentially, which may not help the downstream LLM to answer the question correctly. In this work, we also aim to narrow down the number of correlated tables in a limited ranked list, so that they can best help the LLM to answer the question correctly.
@@ 6 Conclusion
Table retrieval is critical for open-domain question answering, fact-checking, and, more generally, retrieval-augmented generation that relies on table information. However, existing work primarily focuses on single-table retrieval without considering real-life situations where tables are typically normalized. Thus, multiple tables need to be joined to produce sufficient information. Works that consider multi-table retrieval only consider query-table relevance. In this paper, we propose a re-ranking method based on mixed-integer programming to select the best set of tables based on joint decisions about relevance and compatibility. Our experiments demonstrated that our approach outperforms baselines in standard retrieval performance and end-to-end evaluation on the question-answering downstream task.
@@ 7 Limitations
The primary purpose of this paper is to remind the community that table retrieval is still not a solved problem. For complex queries, retrieving multiple (joinable) tables may be required. The proposed MIP-based reranking mechanism in this paper takes a first step to consider join relationships among tables during the inference time. Though MIP may have scalability issues and can be sensitive to the problem data, it provides a flexible framework to model constraints and objectives.
Furthermore, key-foreign-key relationships may not be the only connection among multiple tables for questions in real-life scenarios. Some questions may ask about values of the same column across multiple tables.
We believe exploring a more scalable solution, as well as considering different types of connections among tables in the corpus, would be interesting future work.
Acknowledgments
We thank Mike Cafarella and Çağatay Demiralp for their constructive feedback and discussion on this project. We thank the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the research results reported in this paper. We gratefully acknowledge the support of the DARPA ASKEM project (Award No. HR00112220042), ONR Contract N00014-23-1-2365, and the Croucher Scholarship.